User Tools

Site Tools


handover:crds81

Differences

This shows you the differences between two versions of the page.

Link to this comparison view

Both sides previous revision Previous revision
Next revision
Previous revision
Last revision Both sides next revision
handover:crds81 [2016/01/28 07:23]
Tiege McCarthy
handover:crds81 [2016/01/28 13:38]
Tiege McCarthy
Line 7: Line 7:
      * 17:20 Drivepc restarted in time for experiment (JS)      * 17:20 Drivepc restarted in time for experiment (JS)
      * 17:20 70K stage at 110, 40k stage at 80, pressure at 4.9 will need vacuuming in the morning (JS)      * 17:20 70K stage at 110, 40k stage at 80, pressure at 4.9 will need vacuuming in the morning (JS)
-     * 07:22 UT - Jim reported that the received had been cooled. The stages are currently: ​70K at 79 and 20K at 59. (Ellen)+     * 07:22 UT - Jim reported that the received had been cooled. The stages are currently: 79 at 70K and 59 at 20K. (Ellen) 
 +     * 07:25 UT - Can't connect to newsmerd on monica so I can't check the He supply pressure. (Ellen) 
 +     * 09:39 UT - S/X receiver stages are now at 67 and 53. (Ellen
 +     * 09:40 UT - Tsys values in bands 01 - 08 are approx 110. This is probably due to the current storm. (Ellen) 
 +     * 11:30 UT - Tsys values are back to normal. (Ellen) 
 +     * 13:36 UT - Status monitor reporting date as 1969.293.01.48 UT, schedule appears unaffected. (Bryn)
  
-**Hobart**+**Hobart ​12**
  
 Module: HOB+0075, 6264 GB Module: HOB+0075, 6264 GB
  
      * 17:39 UT - Experiment started OK (JS)      * 17:39 UT - Experiment started OK (JS)
 +     * 09:40 UT - Tsys values all way too high, mostly consistently around 200. This is probably due to the current storm, although Ho system temps are only up to ~110. (Ellen)
 +     * 11:33 UT - X-band has cooled to normal temps but S-band is still too high. (Ellen)
  
 **Katherine** **Katherine**
/home/www/auscope/opswiki/data/pages/handover/crds81.txt · Last modified: 2016/01/28 16:54 by Tiege McCarthy